# MagicDrive: Street View Generation with Diverse 3D Geometry Control

### v1

@@ Abstract
Recent advancements in diffusion models have significantly enhanced the data synthesis with 2D control. Yet, precise 3D control in street view generation, crucial for 3D perception tasks, remains elusive. Specifically, utilizing Bird’s-Eye View (BEV) as the primary condition often leads to challenges in geometry control (e.g., height), affecting the representation of object shapes, occlusion patterns, and road surface elevations, all of which are essential to perception data synthesis, especially for 3D object detection tasks. In this paper, we introduce MagicDrive, a novel street view generation framework offering diverse 3D geometry controls, including camera poses, road maps, and 3D bounding boxes, together with textual descriptions, achieved through tailored encoding strategies. Besides, our design incorporates a cross-view attention module, ensuring consistency across multiple camera views. With MagicDrive, we achieve high-fidelity street-view synthesis that captures nuanced 3D geometry and various scene descriptions, enhancing tasks like BEV segmentation and 3D object detection. 00footnotetext: $*{}^{*}$qual contribution. ${}^{{\dagger}}$orresponding authors. Project Page: https://flymin.github.io/magicdrive.
Figure 1: Multi-camera street view generation from MagicDrive. MagicDrive can generate continuous camera views with controls from road map, object boxes, and text (e.g., weather).
@@ 1 Introduction
The high costs associated with data collection and annotation often impede the effective training of deep learning models. Fortunately, cutting-edge generative models have illustrated that synthetic data can notably boost performance across various tasks, such as object detection (Chen et al., 2023b) and semantic segmentation (Wu et al., 2023). Yet, the prevailing methodologies are largely tailored to 2D contexts, primarily relying on 2D bounding boxes (Lin et al., 2014; Han et al., 2021) or segmentation maps (Zhou et al., 2019) as layout conditions (Chen et al., 2023b; Li et al., 2023).
In autonomous driving applications, a thorough grasp of the 3D environment is essential. This demands reliable techniques for tasks like Bird’s-Eye View (BEV) map segmentation (Zhou & Krähenbühl, 2022; Ji et al., 2023) and 3D object detection (Chen et al., 2020; Huang et al., 2021; Liu et al., 2023; Ge et al., 2023). A genuine 3D geometry representation is crucial for capturing intricate details from 3D annotations, such as road elevations, object heights, and their occlusion patterns, as shown in Figure 2. Consequently, generating multi-camera street-view images according to 3D annotations becomes vital to boost downstream perception tasks.
@@
For street-view data synthesis, two pivotal criteria are realism and controllability. Realism requires that the quality of the synthetic data should align with that of real data; and in a given scene, views from varying camera perspectives should remain consistent with one another (Mildenhall et al., 2020). On the other hand, controllability emphasizes the precision in generating street-view images that adhere to provided conditions: the BEV map, 3D object bounding boxes, and camera poses for views. Beyond these core requirements, effective data augmentation should also grant the flexibility to tweak finer scenario attributes, such as prevailing weather conditions or the time of day. Existing solutions like BEVGen (Swerdlow et al., 2023) approach street view generation by encapsulating all semantics within BEV. Conversely, BEVControl (Yang et al., 2023) starts by projecting 3D coordinates to image views, subsequently using 2D geometric guidance. However, both methods compromise certain geometric dimensions—height is lost in BEVGen and depth in BEVControl.
@@
The rise of diffusion models has significantly pushed the boundaries of controllable image generation quality. Specifically, ControlNet (Zhang et al., 2023a) proposes a flexible framework to incorporate 2D spatial controls based on pre-trained Text-to-Image (T2I) diffusion models (Rombach et al., 2022). However, the challenge of seamlessly integrating multiple 3D conditions with multi-camera view consistency in street view synthesis remains unresolved.
@@
In this paper, we introduce MagicDrive, a novel framework dedicated to street-view synthesis with diverse 3D geometry controls. For realism, we harness the power of pre-trained stable diffusion (Rombach et al., 2022), further fine-tuning it for street view generation. One distinctive component of our framework is the cross-view attention module. This simple yet effective component provides multi-view consistency through interactions between adjacent views. In contrast to previous methods, MagicDrive encodes objects and road maps separately, amplifying the utilization of all available 3D data for controllability enhancement. More specifically, given the sequence-like, variable-length nature of 3D bounding boxes, we employ cross-attention akin to text embeddings for their encoding. Besides, we adapt ControlNet (Zhang et al., 2023a) to view transformation and incorporate road map information. Without resorting to explicit geometric transformations or imposing geometric constraints on multi-camera consistency, our design learns the inherent geometric relationships from data, significantly simplifying our framework. Finally, MagicDrive factors in textual descriptions, offering attribute control such as weather conditions and time of day.
@@
Our MagicDrive framework, despite its simplicity, excels in generating strikingly realistic images that align with road maps, 3D bounding boxes, and varied camera perspectives. Besides, the images produced can enhance the training for both 3D object detection and BEV segmentation tasks. Furthermore, MagicDrive offers comprehensive geometric controls at the scene, background, and foreground levels. This flexibility makes it possible to craft previously unseen street views suitable for simulation purposes. We summarize the main contributions of this work as:
@@
  • The introduction of MagicDrive, an innovative framework that generates multi-perspective camera views conditioned on BEV and 3D data tailored for autonomous driving.
@@
  • The development of simple yet potent strategies to manage 3D geometric data, effectively addressing the challenges of multi-camera view consistency.
@@ 2 Related Work
Diffusion Models for Conditional Generation. Diffusion models (Ho et al., 2020; Song et al., 2020) generate images by learning a progressive denoising process from the Gaussian noise distribution to the image distribution. These models have proven exceptional across diverse tasks, such as text-to-image synthesis (Rombach et al., 2022; Nichol et al., 2022), inpainting (Wang et al., 2023), and instructional image editing (Zhang et al., 2023b; Brooks et al., 2023), due to their adaptability and competence in managing various form of controls (Rombach et al., 2022; Zhang et al., 2023a; Li et al., 2023) and multiple conditions (Liu et al., 2022a; Huang et al., 2023; Gao et al., 2023). Besides, data synthesized from geometric annotations can aid downstream tasks such as 2D object detection (Chen et al., 2023b; Wu et al., 2023). Thus, this paper explores the potential of T2I diffusion models in generating street-view images and benefiting downstream 3D perception models.
Street View Generation. Numerous street view generation models condition on 2D layouts, such as 2D bounding boxes (Li et al., 2023) and semantic segmentation (Wang et al., 2022). These methods leverage 2D layout information corresponding directly to image scale, whereas the 3D information does not possess this property, thereby rendering such methods unsuitable for leveraging 3D information for generation. For street view synthesis with 3D geometry, BEVGen (Swerdlow et al., 2023) is the first to explore. It utilizes a BEV map as a condition for both roads and vehicles. However, the omission of height information limits its application in 3D object detection. BEVControl (Yang et al., 2023) amends the loss of object’s height by the height lifting process, but the projection from 3D to 2D results in the loss of essential 3D geometric information, like depth and occlusion. Thus, neither of them fully exploits 3D annotations or can utilize textual control of driving scenes. In this paper, we propose to encode bounding boxes and road maps separately for more nuanced control and integrate scene descriptions, offering enhanced control over the generation of street views.
Multi-camera Image Generation of a 3D scene fundamentally requires viewpoint consistency. Several studies have addressed this issue within the context of indoor scenes. For instance, MVDiffusion (Tang et al., 2023) employs panoramic images and a cross-view attention module to maintain global consistency, while Tseng et al. (2023) leverage epipolar geometry as a constraining prior. These approaches, however, primarily rely on the continuity of image views, a condition not always met in street views due to limited camera overlap. Our MagicDrive introduces extra cross-view attention modules to UNet, which significantly enhances consistency across multi-camera views.
@@ 4 Street View Generation with 3D Information
The overview of MagicDrive is depicted in Figure 3. Operating on the LDM pipeline, MagicDrive generates street-view images conditioned on both scene annotations ($\mathbf{S}$) and the camera pose ($\mathbf{P}$) for each view. Given the 3D geometric information in scene annotations, projecting all to a BEV map, akin to BEVGen (Swerdlow et al., 2023) or BEVControl (Yang et al., 2023), doesn’t ensure precise guidance for street view generation, as exemplified in Figure 2. Consequently, MagicDrive categorizes conditions into three levels: scene (text and camera pose), foreground (3D bounding boxes), and background (road map); and integrates them separately via cross-attention and an additive encoder branch, detailed in Section 4.1. Additionally, maintaining consistency across different cameras is crucial for synthesizing street views. Thus, we introduce a simple yet effective cross-view attention module in Section 4.2. Lastly, we elucidate our training strategies in Section 4.3, emphasizing Classifier-Free Guidance (CFG) in integrating various conditions.
@@ 4.1 Geometric Conditions Encoding
3D Bounding Box Encoding. Since each driving scene has a variable length of bounding boxes, we inject them through the cross-attention mechanism similar to scene-level information. Specifically, we encode each box into a hidden vector $h^{b}$, which has the same dimensions as that of $h^{t}$. Each 3D bounding box $(c_{i},b_{i})$ contains two types of information: class label $c_{i}$ and box position $b_{i}$. For class labels, we utilize the method similar to Li et al. (2023), where the pooled embeddings of class names ($L_{c_{i}}$) are considered as label embeddings. For box positions $b_{i}\in\mathbb{R}^{8\times 3}$, represented by the coordinates of its 8 corner points, we utilize Fourier embedding to each point and pass through an MLP for encoding, as in Equation 4. Then, we use an MLP to compress both class and position embedding into one hidden vector, as in Equation 5. The final hidden states for all bounding boxes of each scene are represented as $\bm{h}^{b}=[h^{b}_{1}\dots h^{b}_{N}]$, where $N$ is the number of boxes.
@@
Road Map Encoding. The road map has a 2D-grid format. While Zhang et al. (2023a) shows the addictive encoder can incorporate this kind of data for 2D guidance, the inherent perspective differences between the road map’s BEV and the camera’s First-Person View (FPV) create discrepancies. BEVControl (Yang et al., 2023) employs a back-projection to transform from BEV to FPV but complicates the situation with an ill-posed problem. In MagicDrive, we propose that explicit view transformation is unnecessary, as sufficient 3D cues (e.g., height from object boxes and camera pose) allow the addictive encoder to accomplish view transformation. Specifically, we integrate scene-level and 3D bounding box embeddings into the map encoder (see Figure 3). Scene-level embeddings provide camera poses, and box embeddings offer road elevation cues. Additionally, incorporating text descriptions facilitates the generation of roads under varying conditions (e.g., weather and time of day). Thus, the map encoder can synergize with other conditions for generation.
@@ 4.3 Model Training
$\ell=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},t,\{\bm{S},\bm{P}\}}\left[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathcal{E}(\bm{x}_{0})+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t,\{\mathbf{S},\mathbf{P}\})\right].$ (10)
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset (Caesar et al., 2020) as the testing ground for MagicDrive, a prevalent dataset in BEV segmentation and detection for driving. We adhere to the official configuration, utilizing 700 street-view scenes for training and 150 for validation. Our baselines are BEVGen (Swerdlow et al., 2023) and BEVControl (Yang et al., 2023), both recent propositions for street view generation. Our method considers 10 object classes and 8 road classes, surpassing the baseline models in diversity. Appendix B holds additional details.
Evaluation Metrics. We evaluate both realism and controllability for street view generation. Realism is mainly measured using Fréchet Inception Distance (FID), reflecting image synthesis quality. For controllability, MagicDrive is evaluated through two perception tasks: BEV segmentation and 3D object detection, with CVT (Zhou & Krähenbühl, 2022) and BEVFusion (Liu et al., 2023) as perception models, respectively. Both of them are renowned for their performance in each task. Firstly, we generate images aligned with the validation set annotations and use perception models pre-trained with real data to assess image quality and control accuracy. Then, data is generated based on the training set to examine the support for training perception models as data augmentation.
@@
Figure 5: Qualitative comparison with baselines. Both scenes are from nuScenes validation set. We highlight some areas with rectangles to ease comparison. Up: compared with BEVControl, generations from MagicDrive appear more consistent in both background and foreground. Down: compared with BEVControl, image quality of objects from MagicDrive is much better.
Table 2: Comparison about support for 3D object detection model (i.e., BEVFusion). MagicDrive generates 272×736 images for augmentation. Results are reported on the nuScenes validation set. Modality Data mAP $\uparrow$ NDS $\uparrow$ C w/o synthetic data 32.48 37.61 w/ MagicDrive 35.14 +2.66 39.63 +2.02 C+L w/o synthetic data 64.92 69.42 w/ MagicDrive 67.28 +2.36 70.14 +0.72 Table 3: Comparison about support for BEV segmentation model (i.e., CVT). Results are reported by testing on the nuScenes validation set. Data Vehicle mIoU $\uparrow$ Road mIoU $\uparrow$ w/o synthetic data 36.00 74.30 w/ BEVGen 36.60 +0.60 71.90 -2.40 w/ MagicDrive 40.34 +4.34 79.56 +5.26
[Uncaptioned image] Figure 6: Showcase for multi-level control with MagicDrive. We show scene-level, background-level, and foreground-level control separately with diffusion conditions. All scenes are based on the nuScenes validation set.

### v6

@@ Abstract
Recent advancements in diffusion models have significantly enhanced the data synthesis with 2D control. Yet, precise 3D control in street view generation, crucial for 3D perception tasks, remains elusive. Specifically, utilizing Bird’s-Eye View (BEV) as the primary condition often leads to challenges in geometry control (e.g., height), affecting the representation of object shapes, occlusion patterns, and road surface elevations, all of which are essential to perception data synthesis, especially for 3D object detection tasks. In this paper, we introduce MagicDrive, a novel street view generation framework, offering diverse 3D geometry controls including camera poses, road maps, and 3D bounding boxes, together with textual descriptions, achieved through tailored encoding strategies. Besides, our design incorporates a cross-view attention module, ensuring consistency across multiple camera views. With MagicDrive, we achieve high-fidelity street-view image & video synthesis that captures nuanced 3D geometry and various scene descriptions, enhancing tasks like BEV segmentation and 3D object detection. 00footnotetext: $*{}^{*}$qual contribution. ${}^{{\dagger}}$orresponding authors. Project Page: https://flymin.github.io/magicdrive.
Figure 1: Multi-camera street view generation from MagicDrive. MagicDrive can generate continuous camera views with controls from the road map, object boxes, and text (e.g., weather).
@@ 1 Introduction
The high costs associated with data collection and annotation often impede the effective training of deep learning models. Fortunately, cutting-edge generative models have illustrated that synthetic data can notably boost performance across various tasks, such as object detection (Chen et al., 2023c) and semantic segmentation (Wu et al., 2023b). Yet, the prevailing methodologies are largely tailored to 2D contexts, primarily relying on 2D bounding boxes (Lin et al., 2014; Han et al., 2021) or segmentation maps (Zhou et al., 2019) as layout conditions (Chen et al., 2023c; Li et al., 2023b).
In autonomous driving applications, a thorough grasp of the 3D environment is essential. This demands reliable techniques for tasks like Bird’s-Eye View (BEV) map segmentation (Zhou & Krähenbühl, 2022; Ji et al., 2023) and 3D object detection (Chen et al., 2020; Huang et al., 2021; Liu et al., 2023a; Ge et al., 2023). A genuine 3D geometry representation is crucial for capturing intricate details from 3D annotations, such as road elevations, object heights, and their occlusion patterns, as shown in Figure 2. Consequently, generating multi-camera street-view images according to 3D annotations becomes vital to boost downstream perception tasks.
@@
For street-view data synthesis, two pivotal criteria are realism and controllability. Realism requires that the quality of the synthetic data should align with that of real data; and in a given scene, views from varying camera perspectives should remain consistent with one another (Mildenhall et al., 2020). On the other hand, controllability emphasizes the precision in generating street-view images that adhere to provided conditions: the BEV map, 3D object bounding boxes, and camera poses for views. Beyond these core requirements, effective data augmentation should also grant the flexibility to tweak finer scenario attributes, such as prevailing weather conditions or the time of day. Existing solutions like BEVGen (Swerdlow et al., 2023) approach street view generation by encapsulating all semantics within BEV. Conversely, BEVControl (Yang et al., 2023a) starts by projecting 3D coordinates to image views, subsequently using 2D geometric guidance. However, both methods compromise certain geometric dimensions—height is lost in BEVGen and depth in BEVControl.
@@
The rise of diffusion models has significantly pushed the boundaries of controllable image generation quality. Specifically, ControlNet (Zhang et al., 2023a) proposes a flexible framework to incorporate 2D spatial controls based on pre-trained Text-to-Image (T2I) diffusion models (Rombach et al., 2022). However, 3D conditions are distinct from pixel-level conditions or text. The challenge of seamlessly integrating them with multi-camera view consistency in street view synthesis remains.
@@
In this paper, we introduce MagicDrive, a novel framework dedicated to street-view synthesis with diverse 3D geometry controls111 In this paper, our 3D geometry controls contain control from road maps, 3D object boxes, and camera poses. We do not consider others like the exact shape of objects or background contents.. For realism, we harness the power of pre-trained stable diffusion (Rombach et al., 2022), further fine-tuning it for street view generation. One distinctive component of our framework is the cross-view attention module. This simple yet effective component provides multi-view consistency through interactions between adjacent views. In contrast to previous methods, MagicDrive proposes a separate design for objects and road map encoding to improve controllability with 3D data. More specifically, given the sequence-like, variable-length nature of 3D bounding boxes, we employ cross-attention akin to text embeddings for their encoding. Besides, we propose that an addictive encoder branch like ControlNet (Zhang et al., 2023a) can encode maps in BEV and is capable of view transformation. Therefore, our design achieves geometric controls without resorting to any explicit geometric transformations or imposing geometric constraints on multi-camera consistency. Finally, MagicDrive factors in textual descriptions, offering attribute control such as weather conditions and time of day.
@@
Our MagicDrive framework, despite its simplicity, excels in generating strikingly realistic images & videos that align with road maps, 3D bounding boxes, and varied camera perspectives. Besides, the images produced can enhance the training for both 3D object detection and BEV segmentation tasks. Furthermore, MagicDrive offers comprehensive geometric controls at the scene, background, and foreground levels. This flexibility makes it possible to craft previously unseen street views suitable for simulation purposes. We summarize the main contributions of this work as:
@@
  • The introduction of MagicDrive, an innovative framework that generates multi-perspective camera views & videos conditioned on BEV and 3D data tailored for autonomous driving.
@@
  • The development of simple yet potent strategies to manage 3D geometric data, effectively addressing the challenges of multi-camera view consistency in street view generation.
@@ 2 Related Work
Diffusion Models for Conditional Generation. Diffusion models (Ho et al., 2020; Song et al., 2020; Zheng et al., 2023) generate images by learning a progressive denoising process from the Gaussian noise distribution to the image distribution. These models have proven exceptional across diverse tasks, such as text-to-image synthesis (Rombach et al., 2022; Nichol et al., 2022; Yang et al., 2023b), inpainting (Wang et al., 2023a), and instructional image editing (Zhang et al., 2023b; Brooks et al., 2023), due to their adaptability and competence in managing various form of controls (Zhang et al., 2023a; Li et al., 2023b) and multiple conditions (Liu et al., 2022a; Gao et al., 2023). Besides, data synthesized from geometric annotations can aid downstream tasks such as 2D object detection (Chen et al., 2023c; Wu et al., 2023b). Thus, this paper explores the potential of T2I diffusion models in generating street-view images and benefiting downstream 3D perception models.
Street View Generation. Numerous street view generation models condition on 2D layouts, such as 2D bounding boxes (Li et al., 2023b) and semantic segmentation (Wang et al., 2022). These methods leverage 2D layout information corresponding directly to image scale, whereas the 3D information does not possess this property, thereby rendering such methods unsuitable for leveraging 3D information for generation. For street view synthesis with 3D geometry, BEVGen (Swerdlow et al., 2023) is the first to explore. It utilizes a BEV map as a condition for both roads and vehicles. However, the omission of height information limits its application in 3D object detection. BEVControl (Yang et al., 2023a) amends the loss of object’s height by the height-lifting process. Similarly, Wang et al. (2023b) also projects 3D boxes to camera views to guide generation. However, the projection from 3D to 2D results in the loss of essential 3D geometric information, like depth and occlusion. In this paper, we propose to encode bounding boxes and road maps separately for more nuanced control and integrate scene descriptions, offering enhanced control over the generation of street views.
Multi-camera Image Generation of a 3D scene fundamentally requires viewpoint consistency. Several studies have addressed this issue within the context of indoor scenes. For instance, MVDiffusion (Tang et al., 2023) employs panoramic images and a cross-view attention module to maintain global consistency, while Tseng et al. (2023) leverage epipolar geometry as a constraining prior. These approaches, however, primarily rely on the continuity of image views, a condition not always met in street views due to limited camera overlap and different camera configurations (e.g., exposure, intrinsic). Our MagicDrive introduces extra cross-view attention modules to UNet, which significantly enhances consistency across multi-camera views.
@@ 4 Street View Generation with 3D Information
The overview of MagicDrive is depicted in Figure 3. Operating on the LDM pipeline, MagicDrive generates street-view images conditioned on both scene annotations ($\mathbf{S}$) and the camera pose ($\mathbf{P}$) for each view. Given the 3D geometric information in scene annotations, projecting all to a BEV map, akin to BEVGen (Swerdlow et al., 2023) or BEVControl (Yang et al., 2023a), doesn’t ensure precise guidance for street view generation, as exemplified in Figure 2. Consequently, MagicDrive categorizes conditions into three levels: scene (text and camera pose), foreground (3D bounding boxes), and background (road map); and integrates them separately via cross-attention and an additive encoder branch, detailed in Section 4.1. Additionally, maintaining consistency across different cameras is crucial for synthesizing street views. Thus, we introduce a simple yet effective cross-view attention module in Section 4.2. Lastly, we elucidate our training strategies in Section 4.3, emphasizing Classifier-Free Guidance (CFG) in integrating various conditions.
@@ 4.1 Geometric Conditions Encoding
3D Bounding Box Encoding. Since each driving scene has a variable length of bounding boxes, we inject them through the cross-attention mechanism similar to scene-level information. Specifically, we encode each box into a hidden vector $h^{b}$, which has the same dimensions as that of $h^{t}$. Each 3D bounding box $(c_{i},b_{i})$ contains two types of information: class label $c_{i}$ and box position $b_{i}$. For class labels, we utilize the method similar to Li et al. (2023b), where the pooled embeddings of class names ($L_{c_{i}}$) are considered as label embeddings. For box positions $b_{i}\in\mathbb{R}^{8\times 3}$, represented by the coordinates of its 8 corner points, we utilize Fourier embedding to each point and pass through an MLP for encoding, as in Equation 4. Then, we use an MLP to compress both class and position embedding into one hidden vector, as in Equation 5. The final hidden states for all bounding boxes of each scene are represented as $\bm{h}^{b}=[h^{b}_{1}\dots h^{b}_{N}]$, where $N$ is the number of boxes.
@@
Road Map Encoding. The road map has a 2D-grid format. While Zhang et al. (2023a) shows the addictive encoder can incorporate this kind of data for 2D guidance, the inherent perspective differences between the road map’s BEV and the camera’s First-Person View (FPV) create discrepancies. BEVControl (Yang et al., 2023a) employs a back-projection to transform from BEV to FPV but complicates the situation with an ill-posed problem. In MagicDrive, we propose that explicit view transformation is unnecessary, as sufficient 3D cues (e.g., height from object boxes and camera pose) allow the addictive encoder to accomplish view transformation. Specifically, we integrate scene-level and 3D bounding box embeddings into the map encoder (see Figure 3). Scene-level embeddings provide camera poses, and box embeddings offer road elevation cues. Additionally, incorporating text descriptions facilitates the generation of roads under varying conditions (e.g., weather and time of day). Thus, the map encoder can synergize with other conditions for generation.
@@ 4.3 Model Training
$\ell=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},t,\{\bm{S},\bm{P}\}}\left[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathcal{E}(\bm{x}_{0})+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t,\{\mathbf{S},\mathbf{P}\})||\right].$ (10)
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset (Caesar et al., 2020), a prevalent dataset in BEV segmentation and detection for driving, as the testing ground for MagicDrive. We adhere to the official configuration, utilizing 700 street-view scenes for training and 150 for validation. Our baselines are BEVGen (Swerdlow et al., 2023) and BEVControl (Yang et al., 2023a), both recent propositions for street view generation. Our method considers 10 object classes and 8 road classes, surpassing the baseline models in diversity. Appendix B holds additional details.
Evaluation Metrics. We evaluate both realism and controllability for street view generation. Realism is mainly measured using Fréchet Inception Distance (FID), reflecting image synthesis quality. For controllability, MagicDrive is evaluated through two perception tasks: BEV segmentation and 3D object detection, with CVT (Zhou & Krähenbühl, 2022) and BEVFusion (Liu et al., 2023a) as perception models, respectively. Both of them are renowned for their performance in each task. Firstly, we generate images aligned with the validation set annotations and use perception models pre-trained with real data to assess image quality and control accuracy. Then, data is generated based on the training set to examine the support for training perception models as data augmentation.
@@
Figure 5: Qualitative comparison with BEVControl on driving scene from nuScenes validation set. We highlight some areas with rectangles to ease comparison. Compared with BEVControl, generations from MagicDrive appear more consistent in both background and foreground.
Table 2: Comparison about support for 3D object detection model (i.e., BEVFusion). MagicDrive generates 272×736 images for augmentation. Results are reported on the nuScenes validation set. Modality Data mAP $\uparrow$ NDS $\uparrow$ C w/o synthetic data 32.88 37.81 w/ MagicDrive 35.40 +2.52 39.76 +1.95 C+L w/o synthetic data 65.40 69.59 w/ MagicDrive 67.86 +2.46 70.72 +1.13 Table 3: Comparison about support for BEV segmentation model (i.e., CVT). Results are reported by testing on the nuScenes validation set. Data Vehicle mIoU $\uparrow$ Road mIoU $\uparrow$ w/o synthetic data 36.00 74.30 w/ BEVGen 36.60 +0.60 71.90 -2.40 w/ MagicDrive 40.34 +4.34 79.56 +5.26
[Uncaptioned image] Figure 6: Showcase for multi-level control with MagicDrive. We show background-level and foreground-level control separately with different conditions. All scenes are based on the nuScenes validation set. More results are in Appendix E.